Fairfield company applies manufacturing overhead to products at a predetermined rate of $49 per direct labor hour. Its actual manufacturing costs for the most recent period are summarized here:

Direct materials Used on Jobs 101 and 102 $ 89,500
Indirect materials Used on multiple jobs 29,500
Hourly labor wages 750 hours @ $29 per hour
200 hours for Job 101 = $ 5,800
300 hours for Job 102 = 8,700
250 hours for Job 103 = 7,250 21,750
Factory supervision 3,900
Production engineer 7,900
Factory janitorial work 2,400
Selling, general, and administrative salaries 8,900
Other manufacturing overhead costs (factory rent, insurance, depreciation, etc.) 7,400
Other selling, general, and administrative costs (office rent, insurance, depreciation, etc.) 5,900

Manufacturing overhead over applied or under applied

by how much

please show how as well
Answer #1
Direct labor hours             750
Applied Overhead $ 36,750 =750*49
Actual Overhead $ 21,600 =3900+7900+2400+7400
Manfacturing overhead is Overapplied by $15,150 =36750-21600